OpenStudy (phagey):

Suppose you start with at least $52 in your savings account and deposit $27 each week. A. Write an inequality to describe how much money m you have after w weeks. B. If you do this for 11 weeks, will you have enough to buy a bicycle that costs $340? How do you know? For both parts A and B, you must give your answer and explain your answer or show your work.

OpenStudy (anonymous):

A. m=27w+52

OpenStudy (ltrout):

A) 52 + 27w = m. B) 52 + 27w >= 340

OpenStudy (anonymous):

The 27 is how much you make per week and 52 is the original savings.

OpenStudy (anonymous):

B. 340>=27*11+52 340>=349 So yes

OpenStudy (anonymous):

340 is how much you'll need. The greater than or equals to sign is bc you can have more than the required money, the amount of required money, but not less. Times 11 bc it's been 11 weeks.